# Interpretable Plant Leaf Disease Detection Using Attention-Enhanced CNN

### v1

@@ 4.5 Feature Analysis
4.5.1 t-SNE Visualization
To evaluate the discriminative capability of the CBAM-enhanced VGG16 model, we employed t-SNE to project high-dimensional feature representations into a 2-D space. Fig. 11 presents the t-SNE plots on the four grape leaf classes: Esca, Healthy, Leaf Blight, and Black Rot. Each point represents a sample image, colored according to its ground truth class label. It can be interpreted from the t-SNE figure that our proposed model is providing a clear distinguishable cluster separation with minimal class overlap, indicating better feature discrimination. These findings demonstrate how well CBAM works to improve the VGG16 backbone’s representational quality, resulting in improved class-wise separability and model interpretability.
@@
4.5.2 Uniform Manifold Approximation and Projection (UMAP) Feature Visualization
We have also employed UMAP to display features with great dimensions in a two-dimensional area. Fig. 12 presents the UMAP projections of features extracted from the CBAM-VGG16 model. In the CBAM-based UMAP plot, class boundaries are clearly defined, with each category forming dense and distinct clusters. For instance, Healthy samples are sharply isolated from all disease classes, while Black Rot and Esca show minimal overlap, indicating improved intra-class consistency. Leaf Blight forms a distinct cluster far from the rest, highlighting CBAM’s ability to capture unique disease-specific features. These visual differences affirm the effectiveness of CBAM in refining deep feature representations by leveraging spatial and channel-wise attention.
5 Conclusion and Future Work
The study in this work provides an interpretable CBAM-integrated VGG16 approach for detection of plant leaf diseases. By incorporating attention modules at each convolutional stage, the model not only enhances disease-specific feature extraction but also offers improved performance across five public datasets. The performance on five different datasets exhibits the genralizablity of our proposed solution for any other crop. The proposed architecture achieves high classification accuracy while maintaining compatibility with multiple attribution methods for visual explanation. In order to evaluate the model’s interpretability, we carried this in-depth evaluation using different attribution techniques. The qualitative study highlighted that LRP variants produced the most visually clear, localized, and class-discriminative heatmaps with minimal noise. Future directions include refining the attention mechanism to enhance class-awareness and reduce interpretability gaps, exploring global or transformer-inspired attention integration for better contextual representation, and conducting human-in-the-loop evaluations to assess the trust and reliability of visualizations. Furthermore, combining attention maps with attribution heatmaps can lead to a more unified and semantically rich interpretability framework suitable for real-world agricultural issues.
Author Contribution:
@@ 5 Conclusion and Future Work
Funding:
No funding has been received to carry out this work.
Data Availability:

### v2

@@ 4.5 Feature Analysis
To evaluate the discriminative capability of the CBAM-enhanced VGG16 model, we employed t-SNE and UMAP to project high-dimensional feature representations into a 2-D space. Fig. 6 and Fig. 7 presents the t-SNE and UMAP plots on the four grape leaf classes: Esca, Healthy, Leaf Blight, and Black Rot. It can be interpreted from the t-SNE figure that our proposed model is providing a clear distinguishable cluster separation with minimal class overlap, indicating better feature discrimination. In the UMAP plot, class boundaries are clearly defined, with each category forming dense and distinct clusters. These findings demonstrate how well CBAM works to improve the VGG16 backbone’s representational quality, resulting in improved class-wise separability and model interpretability.
@@
5 Conclusion
The study in this work provides an interpretable approach for detection of plant leaf diseases. By incorporating attention modules at each convolutional stage, the model not only enhances disease-specific feature extraction but also offers improved performance across five public datasets. The performance on five different datasets exhibits the genralizability of our proposed solution for any other crop. In order to evaluate the model’s interpretability, we carried this in-depth evaluation using different attribution techniques. The qualitative study highlighted that LRP variants produced the most visually clear, localized, and class-discriminative heatmaps with minimal noise. Future directions include refining the attention mechanism to enhance class-awareness and reduce interpretability gaps, exploring global or transformer-inspired attention integration for better contextual representation, and conducting human-in-the-loop evaluations to assess the trust and reliability of visualisations.
Author Contribution
@@ 5 Conclusion
Funding
The authors declare that no funding was received for this research.
Data Availability